OpenStudy (anonymous):

Find the zeros of f(x) and state the multiplicity of each zero. \[f(x)=x^5+2x^4+x^3\]

OpenStudy (anonymous):

factor out the \(x^3\) first, telling one zero is 0, and it has multiplicity 3

OpenStudy (anonymous):

\[f(x)=x^3(x+1)^2\] Why do we say 0 is a zero of multiplicity 3 is that not contradictory?

OpenStudy (anonymous):

you get \[f(x)=x^3(x^2+2x+1)\] then factor again and get \[f(x)=x^3(x+1)^2\] so your two zeroes are 0 with multiplicity 3 and -1 with multiplicity 2

OpenStudy (anonymous):

"multiplicity" means the exponent on the factored term

OpenStudy (anonymous):

but why would 0 be a zero?

OpenStudy (anonymous):

so for example \((x-2)^2(x+3)^4\) as two zeros, 2 and -3 2 has multiplicity 2 and -3 has multiplicity 4

OpenStudy (anonymous):

because \(f(0)=0\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

each term has an \(x\) in it in fact each term has an \(x^3\) in it

OpenStudy (anonymous):

would it make more sense that the multiplicity is 3 if i wrote \[(x-0)^3(x+1)^2\]?

OpenStudy (anonymous):

maybe that makes it look like all the other examples, but do not write this on a test paper or your teacher will think you are daft. just write \(x^3(x+1)^2\)

OpenStudy (anonymous):

No I could visualize it.

OpenStudy (anonymous):

lol, what's a daft?

OpenStudy (anonymous):

Thanks for all your help!